# A note on Johnson’s rule for minimizing makespan in the Two-Machine Flow Shop scheduling problem

### v1

@@ 3 Computational testing
In flow shop standard instances, the processing times are typically generated according to uniform distributions , see e.g. [8]. Correspondingly, in order to check what is in practice the size of $k_{A}$ and $k_{B}$, we considered instances with sizes $10^{2},10^{3},10^{4}$ and $10^{5}$ and processing times were generated according to uniform distributions $1\dots 10^{2}],[1\dots 10^{3}],[1\dots 10^{4}]$ and $1\dots 10^{5}]$. For each pair (size, distribution), $10$ instances were generated, for a total of 160 instances. The results are summarized in Table $1$. The first column provides the number of jobs. The second column indicates the maximum processing time of the related uniform distribution. The third (fourth) column indicates the maximum value of $k_{A}$ ($k_{B}$) reached among the $10$ generated instances of the same pair (size, distribution). The fifth column indicates for each pair pair (size, distribution) the number of times Property 5 or Property 6 hold. The last column indicates the number of times an optimal sequence can be computed in $O(n)$ time.
From the Table, we evince that, in all instances, both $k_{A}$ and $k_{B}$ where not superior to $6$, meaning that both sets $A$ and $B$ required a limited sorting involving up to $6$ jobs. Correspondingly, for all these instances, the optimal solution of the related $F2\;||\;C_{\max}$ problem could be computed in linear time. Also, for most of the instances ($144/160$) either Property 5 or Property 6 happened to hold, that is just $k_{A}$ or $k_{B}$ jobs needed to be sorted.
njob $p_{max}$ maxmax $k_{A}$ maxmax $k_{B}$ #\# Prop 5,6 #\# $O(n)$
100 100 4 3 5 10
100 1000 3 4 8 10
100 10000 5 4 9 10
100 100000 3 3 7 10
1000 100 5 5 10 10
1000 1000 4 4 10 10
1000 10000 3 4 9 10
1000 100000 3 5 9 10
10000 100 3 6 10 10
10000 1000 4 4 10 10
10000 10000 6 4 8 10
10000 100000 4 6 10 10
100000 100 3 5 10 10
100000 1000 4 4 10 10
100000 10000 5 4 10 10
100000 100000 4 4 9 10
@@ 4 Conclusions
We revisited Johnson’s rule for makespan minimization in the two-machine flow shop problem. We provided a linear time sufficient condition to detect whether the optimum is computable in linear time and showed that, even if instances exist where a complete sorting of subsets $A$ and $B$ is necessary, in practice, on standard uniform distribution, only a very partial sorting is required and therefore an optimal sequence could always be computed in linear time. We remark, also, that all permutations of the jobs in $A$ and $B$ that need not to be sorted can be part of an optimal schedule inducing correspondingly huge subsets of equivalently optimal solutions.

### v3

@@ 4 Computational testing
In this section, we conduct a computational analysis on the sizes of the jobs subsets to be sorted. Specifically, we refer to values $k_{A}$ and $k^{\prime}_{A}$ (with $k^{\prime}_{A}\leqslant k_{A}$) for subset $A$, and to values $\overline{k_{B}}=n-k_{B}-1$ and $\overline{k^{\prime}_{B}}=n-k^{\prime}_{B}-1$ with $(\overline{k^{\prime}_{B}}\leqslant\overline{k_{B}})$ for subset $B$. In flow shop standard instances, the processing times are typically generated according to uniform distributions, see e.g. [11]. Correspondingly, we considered instances with sizes $S=10^{2},10^{3},10^{4}$ and $10^{5}$ where processing times were generated according to uniform distributions $1\dots S]$ and $1\dots 10S]$ for each value of $S$. For each pair size/distribution, $100$ instances were generated, for a total of 800 instances. All algorithms were coded in language C++ 20, and the tests were run on a PC 4.4 GHz AMD Ryzen 7 5800H.
The results are summarized in Table 2. The first column provides the number of jobs. The second column indicates the maximum processing time of the related uniform distribution. The third (fourth) column indicates the maximum value of $k^{\prime}_{A}$ ($k_{A}$) reached among the generated instances of the same pair (size, distribution). The fifth (sixth) column indicates the maximum value of $\overline{k^{\prime}_{B}}$ ($\overline{k_{B}}$) reached among the generated instances of the same pair (size, distribution). The seventh column indicates for each pair (size, distribution) the number of times Property 5 or Property 6 hold. The eighth column indicates the average computational time in seconds of our algorithm. The last column indicates the ratio between the CPU time required by the implementation of Johnson’s algorithm with full sort of sets A and B, and the CPU time required by our approach.
From Table 2, we evince that, in all instances, both $k_{A}$ and $k_{B}$ where not superior to $7$, meaning that both sets $A$ and $B$ required a limited sorting involving up to $7$ jobs. Correspondingly, for all these instances, the optimal solution of the related $F2\;||\;C_{\max}$ problem could be computed in linear time. Also, for most of the instances ($751/800$) either Property 5 or Property 6 happened to hold, that is just $k_{A}$ or $k_{B}$ jobs needed to be sorted.
n $p_{\max}$ $k^{\prime}_{A}$ $k_{A}$ $\overline{k^{\prime}_{B}}$ $\overline{k_{B}}$ Prop5,6 $t_{avg}$ $\tau$
100 100 4 6 3 5 83 $3.0\times 10^{-6}$ 1.57
100 1000 4 5 4 5 80 $3.0\times 10^{-6}$ 1.52
1000 1000 5 6 4 7 96 $1.5\times 10^{-5}$ 8.60
1000 10000 4 5 5 6 96 $1.5\times 10^{-5}$ 8.42
10000 10000 5 7 4 7 98 $1.5\times 10^{-4}$ 30.09
10000 100000 4 5 5 6 99 $1.2\times 10^{-4}$ 38.84
100000 100000 5 6 5 6 100 $1.7\times 10^{-3}$ 56.32
100000 1000000 4 5 5 6 99 $1.9\times 10^{-3}$ 46.06
@@
In addition to the classical uniform distribution on $1,p_{\max}]$, we consider three other discrete probability distributions commonly used to model variability in processing times. For each distribution, parameters are chosen so that the expected processing time is approximately $p_{\max}/2$, ensuring comparability across all families.
  • Negative Binomial Distribution. We use the $(r,p)$ parameterization, where $r>0$ is the number of successes and $0<p<1$ is the probability of success. The mean is $\mathbb{E}[X]=r(1-p)/p$. We fix $r=5$ and choose $p=\frac{r}{r+\mu},\qquad\mu=p_{\max}/2,$ which yields a moderate dispersion while keeping the distribution centered around $p_{\max}/2$.
  • Geometric Distribution. The geometric distribution with parameter $p$ has mean $\mathbb{E}[X]=(1-p)/p$. Setting $p=\frac{2}{p_{\max}+2}$ ensures $\mathbb{E}[X]\approx p_{\max}/2$. This distribution is highly skewed and emphasizes short jobs.
  • Poisson Distribution. The Poisson distribution with rate $\lambda$ has mean $\mathbb{E}[X]=\lambda$. We set $\lambda=p_{\max}/2$, producing a discrete but less heavy-tailed alternative to the geometric and negative binomial laws.
Distribution $k^{\prime}_{A}$ $k_{A}$ $\overline{k^{\prime}_{B}}$ $\overline{k_{B}}$ Prop5,6
geometric 4 6 3 5 96
negative binomial 5 6 4 7 96
poisson 4 7 5 9 95
Table 3: Testing instances with different distribution, where $p_{\max}=1000$ and $n=1000$
Figures 3, 4 and 5 represent the Probability Mass Function (PMF) for these distributions. The results are summarized in Table 3, which shows that the choice of distribution does not impact the values of $k_{A}$ and $k_{B}$ (here the tests were run with $n=1000$ and $p_{\max}=1000$).
Figure 3: Negative Binomial distribution PMF
Figure 4: Geometric distribution PMF
Figure 5: Poisson distribution PMF
@@ 5 Conclusions
We revisited Johnson’s rule for makespan minimization in the two-machine flow shop problem. We provided a sufficient condition, verifiable in linear time, to determine whether the optimum can be computed in linear time. We also showed that, although instances exist for which a complete sorting of subsets $A$ and $B$ is required, in practice, under the standard uniform distribution, only a very limited sorting is needed. As a consequence, an optimal sequence can be computed in linear time with high probability (more than $1-10^{-9}$ for instances with $200$ jobs).
We remark, also, that all permutations of the jobs in $A$ and $B$ that need not to be sorted can be part of an optimal schedule inducing correspondingly huge subsets of equivalently optimal solutions.
Indeed, even if it is out of the scope of this paper, a tailored application of the proposed properties appears suitable for the rapid detection of subsets of distinct optimal sequences, in the spirit of the works of [4] and [6]. For instance, the application of Property 5 to the example proposed in [4] immediately detects $48$ distinct optimal sequences, which is significantly better than the $13$ sequences detected in [4] and corresponds to the value reported in [6]. Further, when a straightforward partial enumeration is combined with a standard application of the proposed properties, $372$ distinct optimal sequences are detected, which is larger than the $324$ optimal sequences mentioned in [6]. An investigation of these issues is most probably worthy of future research.